# AesExpert: Towards Multi-modality Foundation Model for Image Aesthetics Perception

### v1

@@ Abstract.
The highly abstract nature of image aesthetics perception (IAP) poses significant challenge for current multimodal large language models (MLLMs). The lack of human-annotated multi-modality aesthetic data further exacerbates this dilemma, resulting in MLLMs falling short of aesthetics perception capabilities. To address the above challenge, we first introduce a comprehensively annotated Aesthetic Multi-Modality Instruction Tuning (AesMMIT) dataset, which serves as the footstone for building multi-modality aesthetics foundation models. Specifically, to align MLLMs with human aesthetics perception, we construct a corpus-rich aesthetic critique database with 21,904 diverse-sourced images and 88K human natural language feedbacks, which are collected via progressive questions, ranging from coarse-grained aesthetic grades to fine-grained aesthetic descriptions. To ensure that MLLMs can handle diverse queries, we further prompt GPT to refine the aesthetic critiques and assemble the large-scale aesthetic instruction tuning dataset, i.e. AesMMIT, which consists of 409K multi-typed instructions to activate stronger aesthetic capabilities. Based on the AesMMIT database, we fine-tune the open-sourced general foundation models, achieving multi-modality Aesthetic Expert models, dubbed AesExpert. Extensive experiments demonstrate that the proposed AesExpert models deliver significantly better aesthetic perception performances than the state-of-the-art MLLMs, including the most advanced GPT-4V and Gemini-Pro-Vision. Source data will be available at https://github.com/yipoh/AesExpert.
@@ 1. Introduction
Multimodal large language models (MLLMs) have attracted significant attention in the research community (Cai et al., 2023). These foundation models, like GPT-4V (Yang et al., 2023) and LLaVA (Liu et al., 2023b), have demonstrated remarkable progress in serving as general-purpose visual assistants, capable of interacting and collaborating with users (Wu et al., 2024b, 2023a). Concurrently, their outstanding cross-modality capacities further promote a paradigm shift in the computer vision domain, with researchers seeking to transcend the limitations of traditional task-specific approaches and develop multi-modality foundation models that generalize effectively across various visual tasks (Wu et al., 2023b; Liu et al., 2023b; Huang et al., 2020a). Despite the advancements achieved, experiments on current MLLMs reveal obvious limitations in the highly-abstract image aesthetics perception task (Huang et al., 2024b), which covers not only the extensively-studied image aesthetics assessment (IAA) (Yang et al., 2024; Li et al., 2023a), but also fine-grained aesthetic attribute evaluation (e.g., color, light, and composition), aesthetic emotion analysis, and image aesthetics caption (Sheng et al., 2023; Ke et al., 2023; Hentschel et al., 2022). The aesthetic perception abilities of MLLMs are crucial for a wide range of practical applications, such as smart photography, album management, photo recommendation, and image enhancement (Huang et al., 2023, 2020b). Consequently, it is urgent to build a unified foundation model that possesses general abilities across these aesthetic tasks and can precisely respond to open-ended human queries on image aesthetics perception.
@@
Due to the inherently data-hungry nature, current MLLMs rely on vast amounts of multi-modality instruction-following data to meet general-purpose visual and language understanding (Bai et al., 2023). Therefore, researchers have constructed numerous instruction fine-tuning datasets, such as COCO-VQA (Antol et al., 2015), Science QA (Lu et al., 2022) and LLaVA-Instruct-150K (Liu et al., 2023a). However, existing instruction fine-tuning datasets are mainly engineered to enhance the general capacities of MLLMs, e.g. visual question answering (Antol et al., 2015), image captioning (Hossain et al., 2019), object segmentation (Lai et al., 2023) and content understanding (Cai et al., 2023). A notable gap in these datasets is the inadequate focus on visual aesthetics. To address the dilemma, we construct a comprehensively annotated Aesthetic Multi-Modality Instruction Tuning (AesMMIT) dataset, based on which we further fine-tune the open-sourced general foundation models, achieving multi-modality Aesthetic Expert models, dubbed AesExpert, which delivers significantly better aesthetic perception performances than the state-of-the-art MLLMs. Intuitive comparisons and examples are shown in Figure 1. Specifically, this work encompasses three stages:
@@
Stage 1: Collecting human aesthetic feedback from subjective experiments. To bridge the gap between MLLMs and human aesthetics perception, we invite human subjects to provide direct feedback on aesthetic perception and understanding via progressive questions, including three parts: 1) The coarse-grained aesthetic evaluation (e.g. This image looks quite beautiful/unattractive. ). 2) The fine-grained reasoning and explanation based on elemental aesthetic attributes (e.g. clarity, color, light, image object and composition, etc.). 3) The finer-grained description on aesthetic feeling (e.g. novel shooting view, interesting content and expressed emotions). With the three parts, the collected human feedbacks, denoted as AesFeedback, can capture the basic aesthetic perceptions and the evaluation reasoning process. The AesFeedback dataset contains 88K human feedbacks on 21,904 multi-sourced images.
@@
Stage 2: Refining feedback with GPT for instruction-following data. The constructed AesFeedback dataset plays a crucial role in fine-tuning MLLMs for aesthetic instructions. However, to fully harness aesthetic capabilities, the dataset should also encompass an aesthetic question-answering component. To obtain rich question-answer pairs, inspired by the existing works (e.g. COCO-VQA (Lin et al., 2015) and ShareGPT (Chen et al., 2023a)), we leverage GPT to transform human feedback into instruction-following formats, which include both open-end and multiple-choice question-answer pairs. To ensure that MLLMs can handle diverse queries, the instruction-following pairs cover diversified aesthetic perception dimensions (e.g. quality, attribute, emotion, interpretation, enhancement, and context reasoning), and commonly-used question types (e.g. Yes-or-No, What, How, Why, and other open-ended questions). Through the above operations, we obtain the final AesMMIT dataset, which consists of 409K multi-typed instructions to activate stronger aesthetic capabilities.
@@
Stage 3: Building multi-modality aesthetics foundation model based on AesMMIT. We introduce the instruction fine-tuning to improve the open-source MLLMs (Liu et al., 2023a; Ye et al., 2023b) based on the AesMMIT dataset, which not only enables the models to retain their original general knowledge but also facilitates the aesthetics perception capabilities, obtaining the multi-modality Aesthetic Expert models, dubbed AesExpert.
@@
$\bullet$ AesExpert model. We propose multi-modality aesthetic expert models with the aid of the proposed AesMMIT dataset via instruction fine-tuning. Extensive experiments demonstrate that the proposed AesExpert models deliver significantly better aesthetic perception performances than the state-of-the-art MLLMs, including the most advanced GPT-4V and Gemini-Pro.
$\bullet$ Open-source. We release the following sources to the community: (1) the constructed aesthetic multi-modality instruction tuning dataset; (2) the proposed AesExpert models including codes and checkpoints; (3) the visual AesExpert-Chatbot demo. We believe this work would shed light on building more advanced MLLMs with comprehensive aesthetic capabilities.
Figure 2. The proposed dataset construction pipeline. First, we filter 21K diverse-sourced images based on scene, quality and source, and invite 48 human subjects who met the comprehensive criteria. Then, we collect 88K human feedbacks on image aesthetics perception. Finally, human feedbacks are converted into 409K instruction-following pairs (the AesMMIT dataset), which are used for aesthetic instruction tuning.
@@ 2. Related Work
2.1. Multi-modality Foundation Models
@@
Recently, large language models (LLMs), such as GPT-3 (OpenAI, 2023), Flan-T5 (Chung et al., 2022) and LLaMA (Touvron et al., 2023), have made remarkable progress in pure-textual tasks. The success of LLMs has also driven the research on vision-language interaction, resulting in the development of various multi-modality foundation models, e.g. LLaVA (Liu et al., 2023b), MiniGPT-4 (Zhu et al., 2023), mPLUG-Owl (Ye et al., 2023a), Otter (Li et al., 2023b) and Instruct-BLIP (Dai et al., 2023). These models typically contain a pre-trained visual encoder for image processing, an LLM for interpreting instructions and generating responses, and a cross-modality module to align the vision encoder with the LLM (Bai et al., 2023). Despite their impressive performance in general-purpose visual tasks, their performance in handling image aesthetics perception remains underexplored. Therefore, this work is dedicated to building multi-modality aesthetics large language model.
2.2. Multimodal Instruction-following Datasets.
@@
Instruction tuning is an optimization method for multi-modality foundation models, aiming to improve their ability to perform specific tasks (Liu et al., 2023b). With the development of MLLMs, researchers often employ ChatGPT (OpenAI, 2023) or GPT-4 (Yang et al., 2023) to generate diverse and expansive instruction-following data. For example, Zhu et al. (Zhu et al., 2023) employed GPT-3.5 to generate and improve detailed captions for high-quality instruction-following data (MiniGPT-4), which are mainly designed for general visual tasks. Liu et al. (Liu et al., 2023b, a) proposed the multimodal instruction-following dataset (LLaVA-Instruct-150K) based on the existing COCO (Lin et al., 2015) bounding box and caption dataset using GPT-4. In InstructBLIP (Dai et al., 2023), the authors transformed 13 vision-language tasks (e.g. OCR-VQA (Mishra et al., 2019)) into the instruction-following format for instruction tuning. Chen et al. (Chen et al., 2023a) proposed a large-scale image-text dataset featuring 100K highly descriptive captions generated by GPT-4V and 1.2M high-quality captions generated by the proposed caption model, named ShareGPT4V. A more comprehensive survey can be found in (Zhang et al., 2024).
@@
The existing instruction tuning datasets are mainly constructed for general visual tasks, and the lack of image aesthetic data limits the aesthetic perception ability of MLLMs to a large extent. Further, the existing studies suggested that hallucination has become one of the most urgent problems for current MLLMs (Han et al., 2024; Huang et al., 2024a), therefore, using MLLM to directly generate instruction fine-tuning data for aesthetic tasks may further exacerbate this situation. To bridge this gap, this paper presents a corpus-rich aesthetic critique database by collecting human natural language feedback, based on which we establish a comprehensively annotated aesthetic multi-modality instruction tuning dataset, i.e. AesMMIT. Further, we propose multi-modality aesthetic expert models based on aesthetic instruction fine-tuning, achieving significantly better aesthetic perception performances.
@@ 3. Dataset Construction
In this section, we provide a detailed expatiation of the process involved in building the AesMMIT dataset, which is illustrated in Figure 2. Specifically, subsection 3.1 describes the preparation of the subjective experiment including the collection of images and the recruitment of participants. Subsection 3.2 elaborates how we conduct the subjective experiment to obtain the 88K human feedbacks on 21,904 multi-sourced images. Subsection 3.3 explains how we prompt GPT-4 to refine the aesthetic critiques and obtain the 409K instruction-following data covering multiple aesthetic perception dimensions.
@@ 3.1. Experiment Preparation
Image Collection: To guarantee the diversity of image types, we first collect a large number of images from various sources, including natural images (NIs), artistic images (AIs) and artificial intelligence-generated images (AGIs), as shown in Figure 2(a). Then, we use a well-trained scene classification model (Radford et al., 2021) to automatically predict the scene label for each image, based on which we sampled these images to maintain scene diversity. To reduce the long-tail distribution of randomly sampled images (Yang et al., 2024), we add 1239 high-aesthetic images from LITE (Unsplash, 2023) and Impressions (Kruk et al., 2023) datasets, and 1944 low-quality images from SPAQ (Fang et al., 2020) and KonIQ-10K (Hosu et al., 2020) datasets. Finally, a total of 21,904 images are collected, which are further fed to human subjects to collect aesthetic feedbacks based on subjective experiments. Detailed image source of the AesMMIT dataset is summarized in Table 1.
@@
Subject Selection: To ensure the completeness of annotation and the diversity of corpus, we recruit subjects by considering five different perspectives, including age, gender, education, health status and personality, as shown in Figure 2(b). Specifically, we ensure that each subject is in good health and passes the Ishihara color blindness test (Yang et al., 2022). To maintain the quality of annotation, all subjects need to have a high school degree or above, and photography experience. Following the existing works (Yang et al., 2022; Wu et al., 2024a), to ensure the diversity of annotation and language, we further consider age distribution, gender balance, and personality diversity (Li et al., 2020). Finally, considering the difficulty of labeling artistic and artificial intelligence-generated images, we specially invited 6 art practitioners and 9 researchers in the field of image aesthetics evaluation to join the subjective experiment and focus on labeling these two kinds of images. In addition, we also organize two aesthetic knowledge training sessions before the experiment covering evaluation, attribute, emotion and suggestion, aiming to provide them with a more comprehensive understanding of image aesthetics annotation. Finally, based on the principle of voluntary participation (Yang et al., 2022), we invite 48 eligible human subjects to ensure the validity and reliability of annotations.
3.2. Collecting Human Aesthetic Feedback
To enhance the understanding and interpretation capabilities of MLLMs in terms of aesthetics perception, We invite all 48 subjects to contribute their insights and descriptions on image aesthetics through progressive questions, as illustrated in Figure 2(c). This process includes three parts for each test image from coarse to fine. A screenshot of the annotation interface is shown in Figure 3.
2) Fine-grained reasoning and explanation. On the basis of simple aesthetic grade judgments, this question is designed to mine the detailed reasons that influence human aesthetic judgments. Participants need to explain what specific aspects of the image enhance or reduce the aesthetic appeal, such as clarity, color, light, image object and composition. In addition, we encourage participants to provide suggestions for the improvement of low-aesthetic images.
3) Finer-grained description on aesthetic feeling. This question breaks through the description of the inherent aesthetic attributes and aims to explore the impact of human emotions on aesthetic perception when viewing images. Participants need to describe some of the more abstract emotional factors in the image, e.g. novel shooting view, interesting content, expressed emotions, etc.. Furthermore, we encourage participants to analyze the impact of these factors on image aesthetics.
Figure 3. The screenshot of the annotation interface for the subjective experiment.
The setting of the subjective experiment refers to the ITU-R Recommendation BT.500-13 standard (ITU, 2012). By systematically collecting feedback from the above three parts, we construct the corpus-rich aesthetic critique database, called AesFeedback, which contains 88K human feedbacks on 21,904 multi-sourced images. Since these annotations are all provided by human subjects, AesFeedback can capture the characteristics of basic aesthetic perceptions and the evaluation reasoning process, providing a valuable resource for improving the aesthetic capabilities of MLLM.

### v3

@@ Abstract.
The highly abstract nature of image aesthetics perception (IAP) poses significant challenge for current multimodal large language models (MLLMs). The lack of human-annotated multi-modality aesthetic data further exacerbates this dilemma, resulting in MLLMs falling short of aesthetics perception capabilities. To address the above challenge, we first introduce a comprehensively annotated Aesthetic Multi-Modality Instruction Tuning (AesMMIT) dataset, which serves as the footstone for building multi-modality aesthetics foundation models. Specifically, to align MLLMs with human aesthetics perception, we construct a corpus-rich aesthetic critique database with 21,904 diverse-sourced images and 88K human natural language feedbacks, which are collected via progressive questions, ranging from coarse-grained aesthetic grades to fine-grained aesthetic descriptions. To ensure that MLLMs can handle diverse queries, we further prompt GPT to refine the aesthetic critiques and assemble the large-scale aesthetic instruction tuning dataset, i.e. AesMMIT, which consists of 409K multi-typed instructions to activate stronger aesthetic capabilities. Based on the AesMMIT database, we fine-tune the open-sourced general foundation models, achieving multi-modality Aesthetic Expert models, dubbed AesExpert. Extensive experiments demonstrate that the proposed AesExpert models deliver significantly better aesthetic perception performances than the state-of-the-art MLLMs, including the most advanced GPT-4V and Gemini-Pro-Vision. Project homepage: https://yipoh.github.io/aes-expert/.
@@ 1. Introduction
Multimodal large language models (MLLMs) have attracted significant attention in the research community. These foundation models, like GPT-4V (Yang et al., 2023) and LLaVA (Liu et al., 2023b), have demonstrated remarkable progress in serving as general-purpose visual assistants, capable of interacting and collaborating with users (Wu et al., 2023a). Concurrently, their outstanding cross-modality capacities further promote a paradigm shift in the computer vision domain, with researchers seeking to transcend the limitations of traditional task-specific approaches and develop multi-modality foundation models that generalize effectively across various visual tasks (Mi et al., 2023a; Hui et al., 2023; Huang et al., 2020a). Despite the advancements achieved, experiments on current MLLMs reveal obvious limitations in the highly-abstract image aesthetics perception task (Huang et al., 2024c; Zhu et al., 2023c), which covers not only the extensively-studied image aesthetics assessment (IAA) (Zhu et al., 2022, 2023b; Li et al., 2023a), but also fine-grained aesthetic attribute evaluation (e.g. color, light, and composition), aesthetic emotion analysis, and image aesthetics caption (Ke et al., 2023; Hentschel et al., 2022; Huang et al., 2023). The aesthetic perception abilities of MLLMs are crucial for a wide range of practical applications, such as smart photography, album management, photo recommendation, and image enhancement (Huang et al., 2020b; Hui et al., 2022b). Consequently, it is urgent to build a unified foundation model that possesses general abilities across these aesthetic tasks and can precisely respond to open-ended human queries on image aesthetics perception.
@@
Due to the inherently data-hungry nature, current MLLMs rely on vast amounts of multi-modality instruction-following data to meet general-purpose visual and language understanding (Bai et al., 2023). Therefore, researchers have constructed numerous instruction fine-tuning datasets, such as COCO-VQA (Antol et al., 2015), Science QA (Lu et al., 2022) and LLaVA-Instruct-150K (Liu et al., 2023a). However, existing instruction fine-tuning datasets are mainly engineered to enhance the general capacities of MLLMs, e.g. visual question answering (Antol et al., 2015), image captioning (Hossain et al., 2019), object segmentation (Lai et al., 2023) and content understanding (Mi et al., 2023b; Hui et al., 2022a). A notable gap in these datasets is the inadequate focus on visual aesthetics. To address the dilemma, we construct a comprehensively annotated Aesthetic Multi-Modality Instruction Tuning (AesMMIT) dataset, based on which we further fine-tune the open-sourced general foundation models, achieving multi-modality Aesthetic Expert models, dubbed AesExpert, which delivers significantly better aesthetic perception performances than the state-of-the-art MLLMs. Intuitive comparisons and examples are shown in Figure 1. Specifically, this work includes three stages:
@@
Stage 1: Collecting human aesthetic feedback from subjective experiments. To bridge the gap between MLLMs and human aesthetics perception, we invite human subjects to provide direct feedback on aesthetic perception and understanding via progressive questions, including three parts: 1) The coarse-grained aesthetic evaluation (e.g. This image looks quite beautiful/unattractive.). 2) The fine-grained reasoning and explanation based on elemental aesthetic attributes (e.g. clarity, color, light, image object and composition, etc.). 3) The finer-grained description on aesthetic feeling (e.g. novel shooting view, interesting content and expressed emotions). With the three parts, the collected human feedbacks, denoted as AesFeedback, can capture the basic aesthetic perceptions and the evaluation reasoning process. The AesFeedback dataset contains 88K human critiques on 21,904 multi-sourced images.
@@
Stage 2: Refining feedback with GPT for instruction-following data. The constructed AesFeedback dataset plays a crucial role in fine-tuning MLLMs based on aesthetic instructions. However, to fully harness aesthetic capabilities, the dataset should also encompass an aesthetic question-answering component. To obtain rich question-answer pairs, inspired by the existing works (Chen et al., 2023a), we leverage GPT-4 to transform human feedback into instruction-following formats, which include both open-ended and multiple-choice question-answer pairs. To ensure that MLLMs can handle diverse queries, the instruction-following pairs cover diversified aesthetic perception dimensions (e.g. quality, attribute, emotion, interpretation, enhancement, and context reasoning), and commonly-used question types (e.g. Yes-or-No, What, How, Why, and other open-ended questions). Through the above operations, we obtain the final AesMMIT dataset, which consists of 409K multi-typed instructions to activate stronger aesthetic capabilities.
@@
Stage 3: Building multi-modality aesthetics foundation model based on the proposed AesMMIT. We introduce the instruction fine-tuning to improve the open-sourced MLLMs (Liu et al., 2023a; Ye et al., 2023b) based on the AesMMIT dataset, which not only enables the models to retain their original general knowledge but also facilitates the aesthetics perception capabilities, obtaining the multi-modality Aesthetic Expert models, dubbed AesExpert.
@@
$\bullet$ AesExpert model. We propose multi-modality aesthetic expert models with the aid of the proposed AesMMIT dataset via instruction fine-tuning. Extensive experiments demonstrate that the proposed AesExpert models deliver significantly better aesthetic perception performances than the state-of-the-art MLLMs, including the most advanced GPT-4V and Gemini-Pro-Vision.
$\bullet$ Open source. We release the following sources to the community: (1) the constructed aesthetic multi-modality instruction tuning dataset; (2) the proposed AesExpert model, including codes and checkpoints; (3) the visual AesExpert-Chatbot demo. We believe this work would shed light on building more advanced MLLMs with comprehensive aesthetic capabilities.
Figure 2. The proposed dataset construction pipeline. First, we filter 21K diverse-sourced images based on scene, quality and source, and invite 48 human subjects who met the comprehensive criteria. Then, we collect 88K human feedbacks on image aesthetics perception. Finally, human feedbacks are converted into 409K instruction-following pairs (the AesMMIT dataset).
@@ 2. Related Work
2.1. Multi-modality Foundation Model
@@
Recently, large language models (LLMs), such as GPT-3 (OpenAI, 2023), Flan-T5 (Chung et al., 2022) and LLaMA (Touvron et al., 2023), have made remarkable progress in pure-textual tasks. The success of LLMs has also driven the research on vision-language interaction, resulting in the development of various multi-modality foundation models, e.g. LLaVA (Liu et al., 2023b), MiniGPT-4 (Zhu et al., 2023a), mPLUG-Owl (Ye et al., 2023a), Otter (Li et al., 2023b) and Instruct-BLIP (Dai et al., 2023). These models typically contain a pre-trained visual encoder for image processing, an LLM for interpreting instructions and generating responses, and a cross-modality module to align the vision encoder with the LLM (Mi et al., 2024). Despite their impressive performance in general-purpose visual tasks, their performance in handling image aesthetics perception remains underexplored. Therefore, this work is dedicated to building the multi-modality aesthetic large model.
2.2. Multimodal Instruction-following Dataset
@@
Instruction tuning is an optimization method for multi-modality foundation models, aiming to improve their ability to perform specific tasks (Liu et al., 2023b). With the development of MLLMs, researchers often employ ChatGPT (OpenAI, 2023) or GPT-4 (Yang et al., 2023) to generate diverse and expansive instruction-following data. For example, Zhu et al. (Zhu et al., 2023a) employed GPT-3.5 to generate and improve detailed captions for high-quality instruction-following data (MiniGPT-4), which are mainly designed for general visual tasks. Liu et al. (Liu et al., 2023b, a) proposed the multimodal instruction-following dataset (LLaVA-Instruct-150K) based on the existing COCO (Lin et al., 2015) bounding box and caption dataset using GPT-4. In InstructBLIP (Dai et al., 2023), the authors transformed 13 vision-language tasks (e.g. OCR-VQA (Mishra et al., 2019)) into the instruction-following format for instruction tuning. Chen et al. (Chen et al., 2023a) proposed a large-scale image-text dataset featuring 100K highly descriptive captions generated by GPT-4V and 1.2M high-quality captions generated by the proposed caption model, named ShareGPT4V. A more comprehensive survey can be found in (Zhang et al., 2024).
@@
The existing instruction tuning datasets are mainly constructed for general visual tasks, and the lack of image aesthetic data limits the aesthetic perception ability of MLLMs to a large extent. Further, the existing studies suggested that hallucination has become one of the most urgent problems for current MLLMs (Han et al., 2024; Huang et al., 2024b), therefore, using MLLM to directly generate instruction fine-tuning data for aesthetic tasks may further exacerbate this situation. To bridge this gap, this paper presents a corpus-rich aesthetic critique database by collecting human natural language feedback, based on which we establish a comprehensively annotated aesthetic multi-modality instruction tuning dataset, i.e. AesMMIT. Further, we propose multi-modality aesthetic expert models based on aesthetic instruction fine-tuning, achieving significantly better aesthetic perception performances.
@@ 3. Dataset Construction
In this section, we provide a detailed expatiation of the process in building the AesMMIT dataset, which is illustrated in Figure 2. Specifically, subsection 3.1 describes the preparation of the subjective experiment including the collection of images and the recruitment of participants. Subsection 3.2 elaborates how we conduct the subjective experiment to obtain the 88K human feedbacks on 21,904 multi-sourced images. Subsection 3.3 explains how we prompt GPT-4 to refine the aesthetic critiques and obtain the 409K instruction-following data covering multiple aesthetic perception dimensions and commonly-used question types.
@@ 3.1. Experiment Preparation
Image Collection: To guarantee the diversity of image types, we first collect a large number of images from various sources (Xu et al., 2012a), including natural images (NIs), artistic images (AIs) and artificial intelligence-generated images (AGIs), as shown in Figure 2(a). Then, we use a well-trained scene classification model (Radford et al., 2021) to automatically predict the scene label for each image, based on which we sampled these images to maintain scene diversity. To reduce the long-tail distribution of randomly sampled images (Yang et al., 2024; Xu et al., 2011), we add 1239 high-aesthetic images from LITE (Unsplash, 2023) and Impressions (Kruk et al., 2023) datasets, and 1944 low-quality images from SPAQ (Fang et al., 2020) and KonIQ-10K (Hosu et al., 2020) datasets. Finally, a total of 21,904 images are collected, which are further sent to human subjects to collect aesthetic feedbacks based on subjective experiments. Detailed image source of the AesMMIT dataset is summarized in Table 1.
Subject Selection: To ensure the completeness of annotation and the diversity of corpus, we recruit subjects by considering five different perspectives, including age, gender, education, health status and personality, as shown in Figure 2(b). Specifically, we ensure that each subject is in good health and passes the Ishihara color blindness test (Yang et al., 2022). To enhance the quality of annotation, all subjects need to have a high school degree or above, and some experience in photography. Following the existing works (Xu et al., 2013; Wu et al., 2024a; Xu et al., 2012b), to ensure the diversity of annotation, we further consider age distribution, gender balance, and personality diversity (Li et al., 2020). Finally, considering the difficulty of labeling artistic and artificial intelligence-generated images, we specially invited 6 art practitioners and 9 researchers in the field of image aesthetics evaluation to join the subjective experiment and focus on labeling these two kinds of images. In addition, we also organize two aesthetic knowledge training sessions before the experiment covering evaluation, attribute, emotion and suggestion, aiming to provide them with a more comprehensive understanding of image aesthetics annotation. Finally, based on the principle of voluntary participation (Yang et al., 2022), we invite 48 eligible human subjects to ensure the validity and reliability of annotations.
3.2. Collecting Human Aesthetic Feedback
To enhance the understanding and interpretation capabilities of MLLMs in terms of aesthetics perception, we invite 48 subjects to contribute their insights and descriptions on image aesthetics through progressive questions, as illustrated in Figure 2(c). This process includes three parts for each test image from coarse to fine.
@@ 3.2. Collecting Human Aesthetic Feedback
2) Fine-grained reasoning and explanation. On the basis of simple aesthetic grade judgments, this question is designed to mine the detailed reasons that influence human aesthetic judgments (Huang et al., 2024a). In this experiment, all participants need to explain what specific aspects of the image enhance or reduce the aesthetic appeal, such as clarity, color, light, image object and composition. In addition, we encourage participants to provide suggestions for the improvement of low-aesthetics images.
3) Finer-grained description on aesthetic feeling. This question breaks through the description of the inherent aesthetic attributes and aims to explore the impact of human emotions on aesthetic perception when viewing images. Participants need to describe some of the more abstract emotional factors in the image, e.g. novel shooting view, interesting content, expressed emotions, etc. Furthermore, we encourage participants to analyze the impact of these factors on image aesthetics.
The setting of the subjective experiment refers to the ITU-R Recommendation BT.500-13 standard (ITU, 2012). By systematically collecting feedback from the above three parts, we construct the corpus-rich aesthetic critique database, called AesFeedback, which contains 88K human feedbacks on 21,904 multi-sourced images. Since these annotations are all provided by human subjects, AesFeedback can capture the characteristics of basic aesthetic perceptions and the evaluation reasoning process, providing a valuable resource for improving the aesthetic capabilities of MLLMs.
Figure 3. Statistics of top-20 frequently-used words in the AesMMIT database.